# “I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

### v1

@@ A.4. Fairness Metrics and Explanations
A.5. Codebook for Feature Selection Reasons
@@
column Parts refers to Total Number of Participants and Refs refer to Total Number of References Theme Code Description Example Parts. Refs. Fairness and Bias Concerns Affects people who are having fresh start Describes concerns regarding fairness for individuals such as international students or foreign workers who need to establish themselves from scratch in a new system or society. maybe foreign worker, like international students and like people, they have like a starting point. 1 2 Unawareness shouldn’t be a discrimination Emphasizes that individuals unfamiliar with a particular country’s norms or rules due to their background or residency status should not be penalized by AI systems. because I live in a different country. So uh, and I don’t know anything about the loan system here. So so definitely shouldn’t discriminate 3 3 Caring for family cannot be discriminated Points out the necessity for AI systems to accommodate individuals with family caregiving responsibilities without introducing discriminatory bias. For the dependents, I also notice from that graph[data distribution], most of them have one dependent and only a few of them have two dependents. So, that means all of the applicants have dependents and sometimes it does not matter. You need to care for one or two dependents because they’re your family and maybe that’s part of your daily routine and one dependents, two different, doesn’t change too much. So, I think, yeah, because everyone has families so we shouldn’t use that as a reason to reject them 1 1 Concerns about bias Highlights participant concerns regarding biases within AI systems that may unfairly influence decisions affecting people’s lives. if a lot of decisions are falling back on how many dependents, if it’s on the form, how many dependents do you have, then you’re more likely, there will be more weight given to that factor, so they could be denied or accepted and the weight would be too strongly based on dependence, not considering other factors, so I think that’s really one that’s really important to watch out for, because they’re so connected to the other ones 9 11 Doesn’t fully capture the scenario Suggests that some features, such as residence length, may fail to comprehensively represent individuals’ unique situations and therefore require reconsideration in fairness assessments. Residence length cannot approve something. Sometimes people just moved to this new city or country. It cannot help us to fully understand this person’s repayment. One-sided. 7 11 Equality for protected features Stresses the importance of ensuring fairness and equality in AI systems for protected characteristics, particularly to prevent discrimination based on features like residence. And then residence length, So it’s like foreign workers. which means how long they have been living in a country. So yeah, I feel like they shouldn’t judge based on that because that would be discriminating against immigrants. And yeah, it shouldn’t matter how long they have been staying at a place as long as they have a stable job and it’s a stable income. 14 16 Discrimination against Inherent Human Traits and Social Labels Raises issues regarding how factors like residence length might be improperly used to make assessments, potentially creating unfair outcomes for certain groups. I think the first [selected] ones [Age and Gender] are all inherent human traits. Um. Although the second one [foreign worker] doesn’t fit perfectly, they all feel like labels that people naturally carry. If people judge others based on these, it feels a bit like discrimination, like looking at others through tinted glasses. So I rank them higher. 12 17
@@
Table 4 continued from previous page Theme Code Description Example Parts. Refs. Influenced by other features Points out the potential for certain features to be misinterpreted by AI systems, leading to unfair outcomes or decisions. If I have some dependent, I might not own that much, then that will affect my savings. So dependence and savings go together. 2 2 Intention matters Highlights concerns about the potential for AI-driven decisions to inadvertently incorporate bias, particularly in contexts such as residency. The duration is like I think it falls down to the point where whether the intention of the person who is taking this loan is crucial. So whether the duration is longer or shorter should not be discriminated by a AI. 3 3 Not relevant to context-specific decision-making Focuses on the concern that particular features used in AI models might directly lead to discriminatory or unfair outcomes for affected individuals. we shouldn’t really be discriminating based on age, it should be based on whether or not they can pay back the loan […]. 16 22 Financial Ability Considerations Paying long term is fine as long as affordable Emphasizes the need to evaluate individuals’ long-term payment capacity as long as it remains affordable, ensuring AI fairness in financial scenarios. Duration is again just the time that they take to pay the loan back, so just like installment rate, I think it should be fine as long as they can pay it back. So if it makes practical sense, then I don’t think duration should be taken into account as much. 3 4 Avoid prejudging people’s choices Suggests that personal difficulties and hardships should not be considered negative factors when assessing financial ability. Purpose I think […] we cannot prejudge, is that purpose good or bad or is that superior than any other purpose. So, everyone have their own choices. 8 8 Stable income is more necessary Highlights the importance of stable income over other potentially less relevant factors in determining financial capability. dependents also, I don’t see why it’s there. Because, if the person has, okay credit history, the person has a job and can pay and has a salary, has an income, it doesn’t matter if he has children, if he has dependents, or not with their children, or parents, or elderly, like, they’re taking care of. But, I don’t see why they can consider caregivers, for example, not to take loans, or I think it’s also a bit discriminatory. 6 6 Cannot determine repayment capability Points out that certain features might not accurately predict repayment ability and could lead to unfair decisions. Nowadays, people who can earn a lot of money. Part of them are experienced, older business people. Now we have a lot of Internet Celebrity. They are very young. We have many different age ranges. Cannot judge people’s credit level on age 8 8 Don’t capture context-relevant information Acknowledges the complexity of capturing an individual’s complete financial situation through limited features in AI systems. the whole telephone thing, I think that’s just nonsense. Whether someone leaves a phone number or not has just become another data point? 9 16
A.6. Results Table
@@
Custom Metric Detail Consistency & Counterfactual Fairness
Consistency &
Counterfactual Fairness&
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Equalized Odds
@@
Custom Metric Detail
Consistency&
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Counterfactual Fairness &
Conditional Statistical Parity
Consistency &
Conditional Statistical Parity

### v2

@@ A.5. Fairness Metrics and Explanations
A.6. Codebook for Feature Selection Reasons
Table 7 presents the codebook for feature selection reasons.
@@ A.6. Codebook for Feature Selection Reasons
column Parts refers to Total Number of Participants and Refs refer to Total Number of References
Theme Code Description Example Parts. Refs.
Fairness and Bias Concerns Affects people who are having fresh start Describes concerns regarding fairness for individuals such as international students or foreign workers who need to establish themselves from scratch in a new system or society. Maybe foreign worker, like international students and like people, they have like a starting point. 1 2
Unawareness shouldn’t be a discrimination Emphasizes that individuals unfamiliar with a particular country’s norms or rules due to their background or residency status should not be penalized by AI systems. Because I live in a different country. So uh, and I don’t know anything about the loan system here. So so definitely shouldn’t discriminate 3 3
Caring for family cannot be discriminated Points out the necessity for AI systems to accommodate individuals with family caregiving responsibilities without introducing discriminatory bias. For the dependents, I also notice from that graph[data distribution], most of them have one dependent and only a few of them have two dependents. So, that means all of the applicants have dependents and sometimes it does not matter. You need to care for one or two dependents because they’re your family and maybe that’s part of your daily routine and one dependents, two different, doesn’t change too much. So, I think, yeah, because everyone has families so we shouldn’t use that as a reason to reject them 1 1
Concerns about bias Highlights participant concerns regarding biases within AI systems that may unfairly influence decisions affecting people’s lives. If a lot of decisions are falling back on how many dependents, if it’s on the form, how many dependents do you have, then you’re more likely, there will be more weight given to that factor, so they could be denied or accepted and the weight would be too strongly based on dependence, not considering other factors, so I think that’s really one that’s really important to watch out for, because they’re so connected to the other ones 9 11
Doesn’t fully capture the scenario Suggests that some features, such as residence length, may fail to comprehensively represent individuals’ unique situations and therefore require reconsideration in fairness assessments. Residence length cannot approve something. Sometimes people just moved to this new city or country. It cannot help us to fully understand this person’s repayment. One-sided. 7 11
Equality for protected features Stresses the importance of ensuring fairness and equality in AI systems for protected characteristics, particularly to prevent discrimination based on features like residence. And then residence length, So it’s like foreign workers. which means how long they have been living in a country. So yeah, I feel like they shouldn’t judge based on that because that would be discriminating against immigrants. And yeah, it shouldn’t matter how long they have been staying at a place as long as they have a stable job and it’s a stable income. 14 16
@@
Table 7 continued from previous page
Theme Code Description Example Parts. Refs.
Discrimination against Inherent Human Traits and Social Labels Raises issues regarding how factors like residence length might be improperly used to make assessments, potentially creating unfair outcomes for certain groups. I think the first [selected] ones [Age and Gender] are all inherent human traits. Um. Although the second one [foreign worker] doesn’t fit perfectly, they all feel like labels that people naturally carry. If people judge others based on these, it feels a bit like discrimination, like looking at others through tinted glasses. So I rank them higher. 12 17
Influenced by other features Points out the potential for certain features to be misinterpreted by AI systems, leading to unfair outcomes or decisions. If I have some dependent, I might not own that much, then that will affect my savings. So dependence and savings go together. 2 2
Intention matters Highlights concerns about the potential for AI-driven decisions to inadvertently incorporate bias, particularly in contexts such as residency. The duration is like I think it falls down to the point where whether the intention of the person who is taking this loan is crucial. So whether the duration is longer or shorter should not be discriminated by a AI. 3 3
Not relevant to context-specific decision-making Focuses on the concern that particular features used in AI models might directly lead to discriminatory or unfair outcomes for affected individuals. We shouldn’t really be discriminating based on age, it should be based on whether or not they can pay back the loan […]. 16 22
Financial Ability Considerations Paying long term is fine as long as affordable Emphasizes the need to evaluate individuals’ long-term payment capacity as long as it remains affordable, ensuring AI fairness in financial scenarios. Duration is again just the time that they take to pay the loan back, so just like installment rate, I think it should be fine as long as they can pay it back. So if it makes practical sense, then I don’t think duration should be taken into account as much. 3 4
Avoid prejudging people’s choices Suggests that personal difficulties and hardships should not be considered negative factors when assessing financial ability. Purpose I think […] we cannot prejudge, is that purpose good or bad or is that superior than any other purpose. So, everyone have their own choices. 8 8
Stable income is more necessary Highlights the importance of stable income over other potentially less relevant factors in determining financial capability. Dependents also, I don’t see why it’s there. Because, if the person has, okay credit history, the person has a job and can pay and has a salary, has an income, it doesn’t matter if he has children, if he has dependents, or not with their children, or parents, or elderly, like, they’re taking care of. But, I don’t see why they can consider caregivers, for example, not to take loans, or I think it’s also a bit discriminatory. 6 6
Cannot determine repayment capability Points out that certain features might not accurately predict repayment ability and could lead to unfair decisions. Nowadays, people who can earn a lot of money. Part of them are experienced, older business people. Now we have a lot of Internet Celebrity. They are very young. We have many different age ranges. Cannot judge people’s credit level on age 8 8
Don’t capture context-relevant information Acknowledges the complexity of capturing an individual’s complete financial situation through limited features in AI systems. The whole telephone thing, I think that’s just nonsense. Whether someone leaves a phone number or not has just become another data point? 9 16
A.7. Results Table
The table below presents the fairness decision-making results of 26 participants who held outcome fairness perspective.
@@ A.7. Results Table
Custom Metric Detail Consistency & Counterfactual Fairness Consistency & Counterfactual Fairness& Conditional Statistical Parity Consistency & Counterfactual Fairness & Equalized Odds
@@
Custom Metric Detail Consistency& Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Counterfactual Fairness & Conditional Statistical Parity Consistency & Conditional Statistical Parity